A source vibration for a sound wave makes 9.8  106

oscillations (cycles) in 3.5 minutes.
If the speed of the sound wave generated is 318.9 m/s, then determine the distance between
2 successive compressions of the sound wave train.
1
Expert's answer
2020-06-17T09:30:33-0400

The frequency of the source is the number of oscillations in one second:

"f = \dfrac{9.8\cdot 10^6}{3.5\cdot 60s} \approx 46.7 kHz"

The distance between 2 successive compressions of the sound wave is called the wavelength. The wavelength "\lambda", frequency "f" and speed "v" are connected in the following way:


"\lambda = v/f"


Substituting the numerical values, get:


"\lambda = \dfrac{318.9m/s}{46.7\cdot 10^3 Hz} \approx6.81mm"

Answer. 6.81 mm.